OpenStudy (anonymous):

Iris has been studying an invasive population of snails. This particular snail has no local predators, so the population grows wildly. She has observed that the population follows an exponential rate of growth for fifteen years.

OpenStudy (anonymous):

Create your own exponential function, f(x), which models the snail population. You will need to identify the principal population of the snails and the rate of growth each year. Explain to Iris how your function shows the principal population and the rate of growth, in complete sentences. A local snail population grows according to the function g(x) = 200(1.03)2x. Demonstrate the steps to convert g(x) into an equivalent function with only x as the exponent. Then, explain to Iris how the key features of this local snail population compares to the key features of the invasive population. Iris wants to graph the invasive snail population to show the city council. Justify what the appropriate domain and range would be for the function f(x), what the y-intercept would be, and if the function is increasing or decreasing. In five years, a garden festival plans on using the park where Iris has been studying the invasive snails. Explain to the garden festival committee how to find the average rate of change for the snail population between years 2 and 5. Describe what this average rate of change represents.

OpenStudy (anonymous):

@rayne612 @iYuko

OpenStudy (anonymous):

are you having trouble with these or do just dont want to do them...lol

OpenStudy (anonymous):

i dont know how to do them, and i would like for you to guide me through it :)

OpenStudy (anonymous):

you think you can help? :)

OpenStudy (anonymous):

guess what i just found!!!!!!!!!!! :D

OpenStudy (anonymous):

what lol

OpenStudy (anonymous):

SOMEBODY POSTED THE SAME QUESTIONS LOL HOLD ON ILL SHOW U LOL

OpenStudy (anonymous):

Im not an expert at this lol

OpenStudy (anonymous):

Let me try

OpenStudy (anonymous):

I might take a while...but if u want to there is this stie that helps to solve math problems its called Mathway! It is really helpful and it does Algebra too.

OpenStudy (anonymous):

site*

OpenStudy (anonymous):

I use it all the time

OpenStudy (anonymous):

rather have your help though :(

OpenStudy (anonymous):

are my answers correct?

OpenStudy (anonymous):

I dont know i find this one difficult! do u have any note abt how to do this?

OpenStudy (anonymous):

or are u homeschooled

OpenStudy (anonymous):

let me see..

OpenStudy (anonymous):

Exponential functions f(x)=a(b)x, where a is the y-intercept and b is the base of the exponential expression. f(x)=P(1+r)x, where P is the principal amount and r is the rate of change in decimal form. Power of a Power Property A function f(x)=a(b)cx can be represented as f(x)=a(bc)x. Key Features The y-intercept of an exponential function is equal to the a or P, depending on the form used. It is the starting amount in a real-world scenario. The domain for mathematical applications is usually all real numbers. However, limitations could exist depending on the scenario the question is in. The range will have a limit as standard exponential functions do not cross the x-axis. Similar to the domain, careful attention must be paid to the range in real-world scenarios. The average rate of change is the change in the output of the function over a section of the domain. Average Rate of Change = f(b)−f(a) b−a

OpenStudy (anonymous):

Okay now use that to help you. is this homework?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

well an assignment off virtual school

OpenStudy (anonymous):

okay, what dont you understand about how to do this

OpenStudy (anonymous):

well i dont know how to do #3 and i want to know if i did 1 and 2 right

OpenStudy (anonymous):

okay give me one min

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

I have to do some work that is due in 30 mins! I have to get this done! IM SO SORRY

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

aww ok, i understand, good luck with your work!

OpenStudy (anonymous):

jess can you help me? is my answer for 1 and 2 correct?

OpenStudy (anonymous):

im so sorry i feel so guilty!!!! keep tagging that jess girl she will help u

OpenStudy (anonymous):

i mean the last one I tagged

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

dont worry, im sure some one will help me :)

OpenStudy (anonymous):

jess can you help me?